# Advances in 4D Representation: Geometry, Motion, and Interaction

### v1

@@ 2.1.4 Gaussian Splatting based Representations
3D Gaussian Splatting [KKLD23] represents a 3D scene using a set of Gaussian primitives and renders images through rasterization. Each Gaussian is defined by its 3D covariance matrix and spatial position, and is optimized using photometric loss after rasterization. This explicit, discrete representation offers several advantages: it avoids redundant computation in empty regions by only evaluating non-zero Gaussians, enables real-time differentiable rendering through efficient tile-based rasterization [KKLD23], and supports motion modeling more explicitly through deforming each Gaussian. Gaussian Splatting’s explicit formulation and efficient rendering pipeline have made it increasingly favored for dynamic scene applications. Despite these benefits, several challenges remain for this representation in 4D settings.
Temporal Coherency: A straightforward way to model dynamic scene is to generate per-frame 3D Gaussians [RXM24, ZJZ24, SGW24, VNZ25] from an input video. However, maintaining temporal coherence across frames remains difficult due to flickering and blurriness caused by dynamically learned Gaussian colors. To mitigate this, many methods opts for continuous deformation field [YXW23, WGP25, YXL25, XFYX24, SSP24] or tracking-based deformation [SHU24, LWH24] to enforce temporal coherency through local regularization. Another emerging trend is to use native 4DGS that encodes space-time variations within the geometry representation [DWD24, WYF24].
Sparse Input Condition Similar to NeRF, the most comprehensive input for Gassuain Splatting is also multi-view synchronised videos. However, such data scales drastically in volume with video length and number of view angles, along side with the induced computational burden, such data is also harder to acquire. Thus it is a natural choice for the community to shift toward sparse input paradigm. Existing method include simultaneouly learning canonical static Gaussian and deformation field from monocular videos [YPZ24], leveraging additional physical priors such as optical flow [GXC24, LDZY24a], depth [LLW24b] and video diffusion guidance [NCOZMA25, RXM24, WGP25].
Open Challenges. Improving spatial-temporal coherence, training efficiency, and adaptation to sparse inputs remain active research areas. Beyond these, modeling large-scale motion in Gaussian Splatting poses significant challenges, as it requires carefully calibrated covariance matrices to properly re-orient and deform Gaussian primitives across frames. Additionally, while volumetric rendering methods inherently couple appearance and geometry, extracting high-quality explicit meshes from dynamic Gaussian representations remains an important open problem.
@@ 2.2 Structured Representation
Structured representations impose compositional priors that enable more controllable and interpretable 4D modeling. We survey three primary approaches—template-based models, part-based decompositions, and graph-based methods—examining how explicit structural constraints facilitate motion modeling along with the challenges.
@@ 2.2.1 Template-based Representation
Template-based representations combine a parametric mesh with an underlying skeletal structure (kinematic tree). Mesh vertices are bound to the skeleton through a skinning function that determines vertex deformation during skeletal articulation. Animation occurs by manipulating skeletal pose parameters, which drive mesh deformation through skinning weights (detailed in Section˜3.1.1). This representation models object categories sharing common topology and articulated structure—humans, hands, or animals—while allowing individual variation in shape and pose. Examples include SMPL [LMR15] for human bodies, MANO [RTB17] for hands, SMAL [ZKJB17] for animals, FLAME [LBB17] for heads and compositional variants like SMPL+X [PCG19] for body, hand, and face. Templates offer distinct advantages for 4D modeling. They inherently couple motion within parametric structure, provide compactness while maintaining anatomical plausibility, and encode category-specific priors enabling realistic motion and facilitating motion transfer. By decomposing 4D modeling into per-frame pose estimation, templates effectively handle long-term sequences, establishing foundations for digital avatars, human-object interactions, and hand manipulation applications. Despite these advantages, templates face challenges.
@@
Temporal consistency: Per-frame pose estimation can produce temporal jitter from frame-level errors, resulting in unrealistic motion despite correct per-frame geometry. This can be addressed by temporal smoothness constraints through sequential models [KZFM19, KAB20, CMCL21] that leverage temporal context from adjacent frames, or physics-based priors [VXC20] that enforce motion continuity and physical plausibility.
@@ 2.2.2 Part-based Representation
Part-based representation faces two primary challenges: decomposing 3D entities into functional parts and correctly articulating each primitive. The decomposition process requires functional understanding of component relationships and involves both segmentation and completion during part reconstruction. Early approaches rely on segmented datasets [XQM20] but show limited generalizability due to dataset scale constraints. Recent work leverages large language models for semantic reasoning in part segmentation and completion [QYW25, LIC25, XSM25, LLT25]. These methods use LLMs to identify functional parts and plausible articulated motions, then apply amodal completion on segmented 3D parts or retrieval-based reconstruction. An alternative approach learns kinematic part decomposition directly from visual data. SP4D [ZYD25] jointly generates multi-view RGB videos and corresponding kinematic part segmentation from monocular inputs. The method lifts the generated 2D part maps to 3D to derive skeletal structures and harmonic skinning weights, enabling articulated 3D asset creation on extended categories.
Open Challenges. Part-based representation remains a growing field. Active research continues in both 2D [LZRV24a, LZRV24b, VNZ25, SJSC24] and 3D [ZWY24, YQZ24, YGH25, YHZ25], with many techniques yet to be transferred to 4D content creation. Precisely segmenting and articulating parts of 3D models remains an open problem, with current applications concentrated on specific object categories such as furniture and relatively simple motions—prismatic translation, rotation, and revolution. The underlying challenge still lies in functional understanding. Extending to more complex behaviors such as per-part non-rigid deformation and generic objects requires deeper functional understanding of 3D assets and their constituent parts. These directions warrant further investigation by the research community.
@@ 2.2.3 Spatio-Temporal Scene Graphs
Scene graphs are a popular representation for describing multiple geometries in an environment, and the relations between them. 3D scene graphs could be generalized to include a temporal dimension through multiple means. A straightforward approach is to represent a 4D spatio-gemporal scene graphs as a collection of 3D scene graphs at different time stamps [JKFFN20, RCJ21]. While each slice is a standard 3D scene graph, temporal edges can be used to connect the same entity across slices. It is also possible to generalize a linear sequence into an arbitrary graph, where edges link temporally proximal concepts [KGP02]. A collection of slices could also be grouped to represent entire events.
While such representations are effective, they are computationally expensive—when the size and length of the 4D scene grows, because all nodes are still in 3D. A natural generalization will to be adopt 4D nodes, where each node describes not a single 3D geometry, but rather a tube/trajectory of 4D geometries over time [BT11]. In addition to concrete geometries, 4D nodes can also represent abstract concepts such as events. Doing so drastically reduce the size of the graph, removing the need for a node at every time slice. Edges can also be generalized to work with 4D nodes. Instead of distinguishing between spatial (at 3D slices) and temporal (connecting adjatent 3D nodes) edges, a single edge connecting 4D nodes can represent spatio-temporal concepts over time, including time-restricted properties (e.g. co-visibility) [RGA20], actions that evolve through time (e.g. falling from) [YCP23], or causal relationships.
Open challenges: While 4D Spatio-Temporal scene graphs are naturally more suited to describe structural relations and semantics, they are more complex than the unstructured counterparts. Additional effort is needed to curate 4D data that are compatible with such representations. One would also need specific learning algorithms to work with such representations [BT11, YXL18, BPL16]. As a result, it is often harder to achieve the same level of geometric fidelity and temporal granularity, compared to unstructured 4D representations. Connecting such 4D graphs as an additional layer of abstraction over aforementioned unstructured 4D representations is an important future research direction.
@@ 3 Modeling Motion
The core of motion representation is establishing how geometry evolves across temporal frames. This can be formulated as determining the geometric state $\mathcal{G}_{t}$ as time $t$ evolves. We distinguish four principal motion classes in 4D representations: articulated motion, deformation-based motion, tracking-based motion, and hybrid motion (Fig. 3), disscuessed below.
@@ 3.1 Articulated Motion
Articulated motion consists of rigid segments connected by joints that move relative to each other, producing globally non-rigid behavior while each segment undergoes only rigid transformations (translation and rotation). This motion is commonly seen in biological systems (e.g., skeletons) and engineered mechanisms (e.g., robots).
Articulated objects are modeled as kinematic trees [ATK19]-hierarchical structures where joints define degrees of freedom (revolute, prismatic, or fixed) and govern transformation propagation through the hierarchy. Methods for articulated motion divide into two categories. Template-based methods utilize predefined parametric models that encode a known shape and skeletal structure for specific object categories. Template-free methods infers geometry and kinematics directly from observations without category-specific priors, enabling generalization to arbitrary articulated objects at the cost of increased computational complexity and ambiguity in joint inference.
@@ 3.1.1 Template-based Articulation
With many template choices available (see  Section˜2.2.1), we take SMPL [LMR15] as an exemplar for the following discussion, as most template-based methods share similar motion modeling mechanisms. To articulate a standard template mesh from a rest pose to a target pose, SMPL deforms its template mesh using linear blend skinning (LBS), which assigns blend weights to vertices to quantify each joint’s influence. The deformed position $\mathbf{p}_{i}^{m}$ of a canonical point $\mathbf{p}_{i}^{c}$ is computed as:
@@
where $w_{j}(\mathbf{p}_{i}^{c})$ is the blend weight for the $j$-th joint and $\mathbf{T}_{j}\in\mathbb{R}^{4\times 4}$ is its transformation matrix. While LBS works well for surface points, extending it to off-surface points—necessary for volumetric rendering of clothing—is challenging. Traditional methods like barycentric interpolation [JOG24] or nearest-neighbor queries [CYZ24] can produce artifacts with complex deformations, especially for loose garments.
@@
where $\tau(t)$ provides time-varying conditions such as pose parameters. State-of-the-art methods (HumanNeRF [WCS22], TAVA [LTV22]) initialize networks with SMPL weights for improved convergence, then refine through data-driven adjustments. GPS-Gaussian [ZZS24] extends this framework with generalizable part-based neural skinning fields, enabling avatar animation across diverse body shapes from monocular video.
@@ 3.1.2 Template-Free articulation
Joint Parameters: Joint parameters provide a structured representation specifying motion properties including joint type, axis, state, and limits [LXF22]. Common 1-DoF joints (revolute, prismatic) dominate everyday objects, while multi-DoF joints (e.g., ball joints) appear in complex mechanisms. The joint axis encodes orientation and pivot location; joint state represents current configuration (rotation angle or translation distance) [LWY20]. This URDF-standard representation enables direct manipulation [LCAS21] through combined classification (type) and regression (axis, position) [LWY20, GGP25, HIZ23].
@@ 3.1.3 Integrating physical priors
Physical priors ensure plausible reconstructions across both paradigms. Template-based models embed constraints in learned spaces [LMR15], model continuous velocity [SFT25], or learn efficient dynamics [ATFS24]. Collision detection employs spatial partitioning and CCD [RKLM04]. Template-free methods infer consistency from observations, through enforcing SDF and kinematic constraints [WLL25] or incorporating probabilistic limits [SSB11]. Combined constraints, including constant limb lengths [Mal19], connectivity [SSB11], collision-free configurations [WLL25], can also prevent unrealistic motion. Alternatively, integration via neural priors [ATFS24] or differentiable physics [YYj25] enables physically consistent reconstructions for robotics, animation, and simulation.
@@ 3.2 Deformation-guided Representation
This paradigm factorizes dynamic scenes into a static canonical space and time-varying deformation fields [HSY24, PSB21, GCD22, LCM22, ZYX23, RPT23]. The canonical space serves as a reference coordinate system from which all frames derive via learned deformations. For rigid or articulated cases, it corresponds to a neutral pose; for general scenes, it preserves sufficient structure for reliable temporal correspondence.
@@
while backward deformation $\Phi_{\theta}^{-1}$ provides the inverse. Invertibility ensures bidirectional consistency through regularization or invertible architectures [LLW24a, WCC23, CFF22]. NeRF frameworks typically use backward deformation to query canonical attributes [PSB21, PCPMMN21, PSH21], while explicit representations like 3D Gaussians employ forward warping [YGZ24, LLW24a]. For long sequences with substantial motion, multiple local canonical spaces over temporal subwindows maintain coherence while handling large transformations [AHR23, NCOZMA25].
Deformation-based motion representations in 4D tasks encounter key limitations. Topological variations such as object emergence, disappearance, or merging break the bijective mapping assumption, leading to correspondence failures [LNSW21, TTG21]. Large inter-frame motions challenge the smoothness and invertibility of learned deformations, often entangling geometry and appearance [PSH21, PCPMMN21]. Ambiguous correspondences arise under occlusion or low-texture regions where multiple deformations fit equally well [YLSL21, WWXL23] while computational overheads grow rapidly with sequence length and bidirectional consistency [YGZ24, LKLR24]. Lastly, non-rigid or fast motions remain difficult to capture due to temporal discretization limits [PZX21, LSZ22]. Deformation fields can be used to model majorly two types of mottion :(a) rigid and (b) non-rigid motion respectively.

### v2

@@ 2.1.4 Gaussian Splatting based Representations
3D Gaussian Splatting [KKLD23] represents 3D scenes using a set of Gaussian primitives and renders images through rasterization. Each Gaussian is defined by its 3D covariance matrix and spatial position, and is optimized using photometric loss after rasterization. This explicit, discrete representation offers several advantages: it avoids redundant computation in empty regions by only evaluating non-zero Gaussians, enables real-time differentiable rendering through efficient tile-based rasterization [KKLD23], and supports motion modeling more explicitly through deforming each Gaussian. Gaussian Splatting’s explicit formulation and efficient rendering pipeline have made it increasingly favored for dynamic scene applications. Despite these benefits, several challenges remain for this representation in 4D settings.
Temporal Coherency: A straightforward way to model dynamic scenes is to generate per-frame 3D Gaussians [RXM24, VNZ25] from an input video. However, maintaining temporal coherence across frames remains difficult due to flickering and blurriness caused by dynamically learned Gaussian colors. To mitigate this, many methods opt for continuous deformation field [ZJZ24, YXW23, WGP25, YXL25, XFYX24, SSP24] or tracking-based deformation [SHU24, LWH25] or attention-based deformation [SGW24] to enforce temporal coherency through local regularization. Another emerging trend is to use native 4DGS that encodes space-time variations within the geometry representation [DWD24, WYF24].
Sparse Input Conditions: Similar to NeRF, the most comprehensive input for Gaussian Splatting is also multi-view synchronized videos. However, such data scales drastically in volume with video length and number of view angles, along with the induced computational burden, such data is also harder to acquire. Thus it is a natural choice for the community to shift toward a sparse input paradigm. Existing methods include simultaneously learning canonical static Gaussian and deformation field from monocular videos [YPZ24], leveraging additional physical priors such as optical flow [GXC24, LDZY24], depth [LLW25] and video diffusion guidance [NCOZMA25, RXM24, WGP25].
Open Challenges: Improving spatial-temporal coherence, training efficiency, and adaptation to sparse inputs remain active research areas. Beyond these, modeling large-scale motion in Gaussian Splatting poses significant challenges, as it requires carefully calibrated covariance matrices to properly reorient and deform Gaussian primitives across frames. Additionally, while volumetric rendering methods inherently couple appearance and geometry, extracting high-quality explicit meshes from dynamic Gaussian representations remains an important open problem.
@@ 2.2 Structured Representation
Structured representations impose compositional priors that enable more controllable and interpretable 4D modeling. We survey three primary approaches—template-based, part-based, and graph-based methods—examining how explicit structural constraints facilitate motion modeling and the challenges they present.
@@ 2.2.1 Template-based Representation
Template-based representations combine a parametric mesh with an underlying skeletal structure (kinematic tree). Mesh vertices are bound to the skeleton through a skinning function that determines vertex deformation during skeletal articulation. Animation occurs by manipulating skeletal pose parameters, which drive mesh deformation through skinning weights (detailed in Section˜3.1.1). This representation models object categories sharing common topology and articulated structure—humans, hands, or animals—while allowing individual variation in shape and pose. Examples include SMPL [LMR15] for human bodies, MANO [RTB17] for hands, SMAL [ZKJB17] for animals, FLAME [LBB17] for heads, and compositional variants like SMPL+X [PCG19] for body, hand, and face. Templates offer distinct advantages for 4D modeling. They inherently couple motion within parametric structure, provide compactness while maintaining anatomical plausibility, and encode category-specific priors enabling realistic motion and facilitating motion transfer. By decomposing 4D modeling into per-frame pose [GPR23], templates effectively handle long-term sequences, establishing foundations for digital avatars, human-object interactions, and hand manipulation applications. Despite these advantages, templates face challenges.
@@
Temporal Consistency: Per-frame pose estimation can produce temporal jitter from frame-level errors, resulting in unrealistic motion despite correct per-frame geometry. This can be addressed by temporal smoothness constraints through sequential models [KZFM19, KAB20, CMCL21] that leverage temporal context from adjacent frames, or physics-based priors [SGXT20] that enforce motion continuity and physical plausibility.
@@ 2.2.2 Part-based Representation
Part-based representation faces two primary challenges: decomposing 3D entities into functional parts and correctly articulating each primitive. The decomposition process requires functional understanding of component relationships and involves both segmentation and completion during part reconstruction. Early approaches rely on segmented datasets [XQM20] but show limited generalizability due to dataset scale constraints. Recent work leverages large language models for semantic reasoning in part segmentation and completion [QYW25, LIC25, XSM25, LLT25]. These methods use LLMs to identify functional parts and plausible articulated motions, then apply amodal completion on segmented 3D parts or retrieval-based reconstruction. An alternative approach learns kinematic part decomposition directly from visual data. SP4D [ZYD25] jointly generates multi-view RGB videos and corresponding kinematic part segmentation from monocular inputs. The method lifts the generated 2D part maps to 3D to derive skeletal structures and harmonic skinning weights, enabling articulated 3D asset creation for extended categories.
Open Challenges: Part-based representation remains a growing field. Active research continues in both 2D [LZRV24a, LZRV24b, VNZ25, SJSC24, WGPYZ24] and 3D [ZWY24, YQZ24, YGH25, YHZ25], with many techniques yet to be transferred to 4D content creation. Precisely segmenting and articulating parts of 3D models remains an open problem, with current applications concentrated on specific object categories such as furniture and relatively simple motions—prismatic translation, rotation, and revolution. The underlying challenge still lies in functional understanding. Extending to more complex behaviors such as per-part non-rigid deformation and generic objects requires a deeper functional understanding of 3D assets and their constituent parts. These directions warrant further investigation by the research community.
@@ 2.2.3 Spatio-Temporal Scene Graphs
Scene graphs are a popular representation for describing multiple geometries in an environment and the relations between them. 3D scene graphs could be generalized to include a temporal dimension through multiple means. A straightforward approach is to represent a 4D spatio-temporal scene graph as a collection of 3D scene graphs at different time stamps [JKFFN20, RCJ21]. While each slice is a standard 3D scene graph, temporal edges can be used to connect the same entity across slices. It is also possible to generalize a linear sequence into an arbitrary graph, where edges link temporally proximal concepts [KGP02]. A collection of slices could also be grouped to represent entire events.
While such representations are effective, they are computationally expensive—when the size and length of the 4D scene grow, because all nodes are still in 3D. A natural generalization is to adopt 4D nodes, where each node describes a tube/trajectory of 4D geometries over time [BT11], rather than a single 3D geometry. In addition to concrete geometries, 4D nodes can also represent abstract concepts such as events. Doing so drastically reduces the size of the graph, removing the need for a node at every time slice. Edges can also be generalized to work with 4D nodes. Instead of distinguishing between spatial (at 3D slices) and temporal (connecting adjacent 3D nodes) edges, a single edge connecting 4D nodes can represent spatio-temporal concepts over time, including time-restricted properties (e.g. co-visibility) [RGA20], evolving actions (e.g. falling from) [YCP23], or causal relationships.
Open Challenges: While 4D Spatio-Temporal scene graphs naturally describe structural relations and semantics better, they are more complex than the unstructured counterparts. Curating compatible 4D data and developing specific learning algorithms for such representation also require significant efforts  [BT11, YXL18, BPL16]. Consequently, achieving comparable geometric fidelity and temporal granularity to unstructured 4D representations remains challenging. Connecting such 4D graphs as an additional layer of abstraction over aforementioned unstructured 4D representations is an important future research direction.
@@ 3 Modeling Motion
The core of motion representation is establishing how geometry evolves across temporal frames. This can be formulated as determining the geometric state $\mathcal{G}_{t}$ as time $t$ evolves. We distinguish four principal motion classes in 4D representations: articulated motion, deformation-based motion, tracking-based motion, and hybrid motion (Fig. 3), discussed below.
@@ 3.1 Articulated Motion
Articulated motion consists of rigid segments connected by joints that move relative to each other, producing globally non-rigid behavior while each segment undergoes only rigid transformations (translation and rotation). This motion is commonly seen in biological systems (e.g., skeletons) and engineered mechanisms.
Articulated objects are modeled as kinematic trees [ATK19]-hierarchical structures where joints define degrees of freedom (revolute, prismatic, or fixed) and govern transformation propagation through the hierarchy. Methods for articulated motion fall into two categories. Template-based methods utilize predefined parametric models that encode a known shape and/or skeletal structure for specific object categories. Template-free methods infer geometry and kinematics directly from observations without category-specific priors, enabling generalization to arbitrary articulated objects at the cost of increased computational complexity and ambiguity in joint inference.
@@ 3.1.1 Template-based Articulation
Among available template choices (see  Section˜2.2.1), we take SMPL [LMR15] as an exemplar for the following discussion, as most template-based methods share similar motion modeling mechanisms. To articulate a standard template mesh from a rest pose to a target pose, SMPL deforms its template mesh using linear blend skinning (LBS), which assigns blend weights to vertices to quantify each joint’s influence. The deformed position $\mathbf{p}_{i}^{m}$ of a canonical point $\mathbf{p}_{i}^{c}$ is computed as:
@@
where $w_{j}(\mathbf{p}_{i}^{c})$ is the blend weight for the $j$-th joint and $\mathbf{T}_{j}\in\mathbb{R}^{4\times 4}$ is its transformation matrix. This provides natural support for in-domain applications such as human motion synthesis [YSI23] and reconstruction [GPR23], where tasks are accomplished through template parameter prediction. Further combination with visually rich representations such as Gaussian Splatting [LZWL24] allows realistic motion and appearance to be generated simultaneously.
@@
where $\tau(t)$ provides time-varying conditions such as pose parameters. Instead of using the fully equipped SMPL template, this branch of work, at minimum, takes only the skeleton as input and derives the skinning weight. HumanNeRF [WCS22], TAVA [LTV22], and A-NeRF [PDW21] employ off-the-shelf modules to predict posed skeletons as the prior and model motion sequence from input videos. LS-Avatar [SWS25] applies part-based neural skinning fields, enabling realistic avatar animation from monocular video. NeuroSkinning [LZT19] and SNARF [CZB21] take the skeletons as input and infer skinning weights to synthesize motion, with the former demonstrating applicability on cloth animations once the skeleton is provided.
@@ 3.1.2 Template-Free articulation
Joint Parameters: Joint parameters provide a structured representation specifying motion properties including joint type, axis, state, and limits [LXF22]. Common 1-DoF joints (revolute, prismatic) dominate everyday objects, while multi-DoF joints (e.g., ball joints) appear in complex mechanisms. The joint axis encodes orientation and pivot location; joint state represents current configuration (rotation angle or translation distance) [CSWK24].
@@ 3.1.3 Integrating physical priors
Physical priors ensure plausible reconstructions across both paradigms. Template-based models embed constraints in learned spaces [LMR15], model continuous velocity [SFT25], or learn efficient dynamics [ATFS24]. Collision detection employs spatial partitioning and CCD [RKLM04]. Template-free methods infer consistency from observations, through enforcing SDF and kinematic constraints [WLL25] or incorporating probabilistic limits [SSB11]. Alternatively, integration via neural priors [ATFS24] or physics constraint injection [YSI23] enables physically consistent reconstructions for robotics, animation, and simulation.
@@ 3.2 Deformation-guided Representation
This paradigm factorizes dynamic scenes into a static canonical space and time-varying deformation fields [HSY24, PSB21, GCD22, LCM22, ZYX23, RPT23]. The canonical space serves as a reference from which all frames derive via learned deformations. For rigid or articulated cases, it corresponds to a neutral pose; for general scenes, it preserves sufficient structure for reliable temporal correspondence.
@@
while backward deformation $\Phi_{\theta}^{-1}$ provides the inverse and $\tau(t)$ is a function of time. Invertibility ensures bidirectional consistency through regularization or invertible architectures [LLW25, WCC23, CFF22]. NeRF frameworks typically use backward deformation to query canonical attributes [PSB21, PCPMMN21, PSH21], while explicit representations like 3D Gaussians employ forward warping [YGZ24, LLW25]. For long sequences with substantial motion, multiple local canonical spaces over temporal subwindows maintain coherence while handling large transformations [AHR23, NCOZMA25].
Deformation-based motion representations in 4D tasks encounter key limitations. Topological variations such as object emergence, disappearance, or merging break the bijective mapping assumption, leading to correspondence failures [LNSW21, TTG21]. Large inter-frame motions challenge the smoothness and invertibility of learned deformations, often entangling geometry and appearance [PSH21, PCPMMN21]. Ambiguous correspondences arise under occlusion or low-texture regions where multiple deformations fit equally well [YLSL21, WWXL23] while computational overheads grow rapidly with sequence length and bidirectional consistency [YGZ24, LKLR24]. Lastly, non-rigid or fast motions remain difficult to capture due to temporal discretization limits [PZX21, LSZ22]. Deformation fields model two main types of motion: (a) rigid and (b) non-rigid motion.